OpenStudy (anonymous):

2. At the baseball concessions, 4 hotdogs and 2 sodas cost $8. The cost of 5 hotdogs and 3 drinks is $10.50. Determine the cost of a hotdog and the cost of a soda. 3. Two angles are supplementary if the sum of their measures is 180 degrees. If the larger of two supplementary angles is 30 degrees more than the smaller angle, find the measures of the two angles. 4. According to the Guinness Book of World Records, the sum of the heights of the shortest and tallest persons on record is 130 inches. If the difference in their heights was 84 inches, how tall was each? 5. Mark earns a weekly

OpenStudy (anonymous):

2) hotdog=X, soda=Y 4X + 2Y = 8---------- x5 5X + 3Y = 10.50------ x4 ================= 50X+10Y=40 20X+12Y=42 ______________- 2Y=2 Y=$1.... can u continue? :)